OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -6, an = 4 • an-1

OpenStudy (anonymous):

You are given with: \[a_n = 4 \cdot a_{n-1}\] Put n = 2 here and find \(a_2\), can you do it??

OpenStudy (anonymous):

@waterineyes a(2)=4 x a(2)-1 : a2^2=-4?

OpenStudy (anonymous):

-4??

OpenStudy (anonymous):

Do not mention me, I am here only..

OpenStudy (anonymous):

What you got for \(a_2\) ??

OpenStudy (anonymous):

@Ambition I am asking you only...

OpenStudy (anonymous):

@Ambition if you will not reply, then you will not get it..

OpenStudy (anonymous):

I got -4

OpenStudy (anonymous):

How?? Can you show me in a nice step??